Student tuition at Boehring University is $150 per semester credit hour. The state supplements school revenue by $90 per semester credit hour. Average class size for a typical 3-credit course is 55 students. Labor costs are $5,200 per class, materials costs are $23 per student per class, and overhead costs are $27,000 per class. The multifactor productivity ratio currently is 1.18 and the labor productivity ratio is $176.79 per hour if the instructors work on an average of 14 hours per week for 16 weeks for each 3-credit class of 55 students Coach Bjourn Toulouse led the Big Red Herrings to several disappointing football seasons. Only better recruiting will return the Big Red Herrings to winning form Because of the current state of the program, Boehring University fans are unlikely to support increases in the $192 season ticket price. Improved recruitment will increase overhead costs to $33,000 per class section from the current $27,000 per class section. The universitys budget plan is to cover recruitment costs by increasing the average class size to 75 students. Labor costs will increase to $6,000 per 3-credit course. Material costs will be about $30 per student for each 3-credit course. Tuition will be $200 per semester credit, which is supplemented by state support of $90 per semester credit a. The multifactor productivity ratio with the universitys plan to meet the expenses related to improving recruitment is 1.58. (Enter your response rounded to two decimal places.) As compared to the present multifactor productivity ratio, the universitys plan will result in an improved multifactor productivity ratio b. If instructors are going to work an average of 20 hours per week for 16 weeks for each 3-credit class of 75 students, the labor productivity ratio is $ per hour. (Enter your response rounded to two decimal places.) I need help on part b
Answer #1

1. Scenario I- Income = ( 150+90) x14x16 x55 = $2956800

Expenses = 5200+27000+23x55 = 33465

Net income = 2956800-33465 =2923335

Labour hours put = 16x14 =224

Labour productivity = 2923335/224 = 13050.60

Number of insructors = Productivity / productivity ratio = 13050.6/ 176.79 =74

LPR in case II

Income = 290x75 x20x16 =6960000

Expenses = 6000+30x75+33000 = 41250

Productivity per hour = 6960000-41250 / 320 = 21621.09

Productivity ratio = 21621.09 / 74 = 292.17